# 3D Foundation AI Model for Generalizable Disease Detection in Head Computed Tomography

### v1

@@ Abstract
Head computed tomography (CT) imaging is a widely-used imaging modality with multitudes of medical indications, particularly in assessing pathology of the brain, skull, and cerebrovascular system. It is commonly the first-line imaging in neurologic emergencies given its rapidity of image acquisition, safety, cost, and ubiquity. Deep learning models may facilitate detection of a wide range of diseases. However, the scarcity of high-quality labels and annotations, particularly among less common conditions, significantly hinders the development of powerful models. To address this challenge, we introduce FM-CT: a Foundation Model for Head CT for generalizable disease detection, trained using self-supervised learning. Our approach pre-trains a deep learning model on a large, diverse dataset of 361,663 non-contrast 3D head CT scans without the need for manual annotations, enabling the model to learn robust, generalizable features. To investigate the potential of self-supervised learning in head CT, we employed both discrimination with self-distillation and masked image modeling, and we construct our model in 3D rather than at the slice level (2D) to exploit the structure of head CT scans more comprehensively and efficiently. The pre-training phase is followed by fine-tuning on smaller, annotated downstream datasets, thereby optimizing the model for specific diagnostic tasks, such as detecting hemorrhages, tumors, and other abnormalities. The model’s downstream classification performance is evaluated using internal and three external datasets, encompassing both in-distribution (ID) and out-of-distribution (OOD) data. Our results demonstrate that the self-supervised foundation model significantly improves performance on downstream diagnostic tasks compared to models trained from scratch and previous 3D CT foundation models on scarce annotated datasets. Furthermore, the model maintains strong generalization across different datasets, indicating its potential for broad clinical applicability. This work highlights the effectiveness of self-supervised learning in medical imaging and sets a new benchmark for head CT image analysis in 3D, enabling broader use of artificial intelligence for head CT-based diagnosis.
@@ Introduction
Head computed tomography (CT) is often the first step in diagnosing a wide range of neurological disorders, including head trauma, hemorrhages, hydrocephalus, and malignancies. Head CT scans are faster, more accessible, and generally less expensive than magnetic resonance imaging (MRIs), making them ideal for emergencies like traumatic brain injury (TBI) or suspected stroke. They are also effective in detecting bone fractures, or neurovascular pathologies such as arterial venous malformations. Despite its widespread use, CT lacks the contrast resolution and hence the sensitivity for many disorders dependent upon diagnosis by MRI, thus MRI is the imaging modality of choice for many neurologic diseases. MRI, however, is more costly, risks potential heating or displacement of indwelling implants, and suffers generally slower acquisition times, increasing patient discomfort and risking non-diagnostic examinations due to its greater sensitivity to motion-related artifacts. It is also more expensive than CT and is contraindicated in specific patients. Access to MRI is a major challenge in resource-limited countries. The timely and arduous determination of certain pathologies can delay appropriate medical and surgical treatment for patients. There is significant potential to harness artificial intelligence (AI) algorithms to enhance the diagnostic and early detection capabilities of head CT, providing critical support in clinical decision-making and improving patient outcomes. Early and accurate diagnosis can potentially lead to more effective treatments, reduce complications, and improve patient survival.
@@
Current research on AI-driven diagnosis using head CT is limited due to both lack of data availability and the complexity of model architectures. Although datasets such as RSNA [1] and CQ500 [2] provide publicly available head CT data, they remain small (RSNA includes approximately 10K samples and CQ500 approximately 1K), and their primary focus is on hemorrhage detection, which restricts broader applicability as a credible path to clinical decision support. Moreover, many existing models are designed with highly task-specific architectures that may not generalize well to diverse clinical applications. These models typically apply 2D convolution neural networks (CNN) to sequentially process 3D volumes slice-by-slice under the supervision of slice-level labels [3, 2, 4]. Slice-level labels are often expensive to acquire, and models trained on 2D slices often struggle to generalize to conditions like neurodegenerative diseases, where slice-level labels are not easily defined. Developing models that can harness the information embedded within the 3D structure of CT images while requiring minimal slice-level labeled data can thus expand the impact of such approaches. To address these challenges, we have developed FM-CT: a Foundation Model for Head CT, and demonstrated robust performance across multiple tasks and datasets, which highlights our model’s potential for broad clinical applicability.
@@
In this work, we introduce FM-CT, a 3D foundation model for head CT scans, developed using self-supervised learning (SSL). While SSL has shown success in natural images [21, 22, 23, 24, 25, 26, 6, 7, 27, 28] and in medical imaging [29, 30, 12, 14, 31, 32, 13, 33, 34, 35, 36], training a robust 3D CT volume encoder presents distinct challenges, such as selecting appropriate pretext tasks, managing spatial normalization, and addressing high computational demands. To overcome these challenges, we developed a standardized pipeline that normalizes head CT scans from various protocols, producing consistent input for our foundation model. For pre-training, we adapted two SSL frameworks—self-distillation, inspired by DINO [24], and masked prediction, inspired by MAE [26]. These methods were tailored to train a volumetric encoder based on a customized vision Transformer (ViT)[37]. The full details of our design and choices are described in Section“Method”. pre-training was conducted on a large-scale dataset comprising 361,663 head CT scans from a major clinical institution.
@@
To evaluate the foundation model, we systematically assessed its performance and generalizability across 10 downstream disease detection tasks using diverse internal and external datasets, as illustrated in Fig. 1. Beyond commonly studied hemorrhages, our evaluation includes crucial yet less-explored tasks in head CT, such as identifying brain tumors, Alzheimer’s disease and related dementia (ADRD), edema, and hydrocephalus (HCP). For each downstream task, the foundation model was fine-tuned using task-specific labels. Given the scarcity of expert-annotated public datasets for these conditions, we leverage electronic health records (EHR) to acquire labels of each task. While EHRs may include missing data and suffer potential label-noise, they remain a valuable and practical source for large-scale patient status labeling that can be used to evaluate the performance of the foundation model.
@@
Our results reveal substantial performance improvements enabled by our foundation model. Downstream models initialized with the pre-trained weights of foundation model achieve a $16.07\%$ improvement in macro-AUC over models trained from scratch with random initialization on internal NYU Langone data, and $20.86\%$ and $12.01\%$ improvements on external datasets from NYU Long Island (previously a separate hospital) and RSNA, respectively ($P<0.001$ for all comparisons). These findings underscore the potential of our foundation model to advance AI-based interpretation of head CT scans, supporting more accurate diagnosis and early disease detection. Furthermore, as described in Section “Results”, we demonstrate the model’s capabilities in out-of-distribution generalization (Fig. 1), few-shot learning (Fig. 2), and scalability (Fig. 3), highlighting the method’s potential in scenarios with limited annotated fine-tuning data, or scenarios such as federated learning which provide access to orders of magnitude more data. Overall, the experimental results on multiple datasets and tasks underscore the generalizability, adaptability and effectiveness of the model, and pave the way for significant impact in real-world clinical applications.
@@
Figure 1: Overview of the study - the approach to developing a foundation model for head CT and its performance in disease detection tasks. a, Collection of training data and pretraining of the foundation model. b, Query disease labels associated with head CT scans for downstream tasks. c, Evaluation design of the foundation model using both internal and external datasets. d, Application of the foundation model to various downstream disease detection tasks. e-g, The performance comparison among different scenarios: (e) training with and without foundation model, (f) comparison of our CT foundation model vs. other CT foundation models, (g) in-domain finetuning and external validation in transfer learning.
@@ Results
The key aim of the foundation model is to develop a single model that improves performance on a wide range of downstream tasks of detecting recognizable abnormalities from head CT scans. To evaluate the capability of the foundation model, we train classification models for multiple disease detection tasks by fine-tuning the foundation model separately per disease, and assessing the fine-tuned model’s performance on held-out validation and external data sets. The selected downstream tasks include detecting various types of hemorrhages (intraparenchymal hemorrhage (IPH), intraventricular hemorrhage (IVH), subdural hemorrhage (SDH), epidural hemorrhage (EDH), subarachnoid hemorrhage (SAH), and intracranial hemorrhage (ICH)), brain tumors, hydrocephalus (HCP), edema, and Alzheimer’s diseases and related dementia (ADRD). Fig. 1a,b,c show the overview of our pre-training framework and included data, EHR-matching, and datasets used in pre-training, in-domain fine-tuning, and external validation. Overall N=361,663 scans were used during pre-training, and four distinct datasets from different sources were used for various forms of validation (NYU Langone N=26,487; NYU Long Island N=2,202; RSNA N=1,058; and CQ500 N=236). NYU Langone is a hospital system comprised of multiple geographically distinct hospitals including two Level 1 Trauma Centers and three Comprehensive Stroke Centers. NYU Long Island, a Level 1 Trauma Center/Comprehensive Stroke Center, is treated as an external dataset for the purposes of this study.
@@
The first two rows of Fig. 1e report the task-specific AUCs for Vision Transformer (ViT) classifiers trained from scratch with random initialization, namely scratch, versus those fine-tuned from the foundation model, namely fine-tuned on NYU Langone data. The fine-tuned models consistently outperform the scratch model across all 10 disease detection tasks, achieving a macro-AUC of 0.852 — a $16.07\%$ increase over the scratch model’s 0.734 ($P<0.001$). Additionally, in Fig. 1f and Supplementary Fig. 5 we compared the foundation model with two other foundation model for 3D CT scans — Merlin [20] and Google’s CT Foundation[38] model. Merlin outperforms the scratch model with a macro-AUC of $5.67\%$ while falling short compared to our foundation model with $7.51\%$ lower macro-AUC ($P<0.001$, illustrated in Supplementary Fig. 5). Although Merlin is not directly comparable to our foundation model as it was pre-trained on abdominal CT, it still provides a valuable baseline. We compare our model to Google CT Foundation model with linear probing, because trainable weights for end-to-end fine-tuning are not provided for this model. We consistently observe improved model performance across the board (in Fig. 1f and Supplementary Fig. 7). These findings demonstrate that despite the progress in general domain multimodal models, specialized foundation model pre-trained on head CT data still significantly enhance the understanding of brain CTs.
@@
To assess our foundation model’s generalization to out-of-distribution data, we compiled three external datasets from multiple institutions and sources: NYU Long Island, RSNA [1], and CQ500 [2], as shown in Fig. 1c (NYU Langone and NYU Long Island are geographically separate and distinct institutions within the broader health system). The data in these external datasets has a different distribution than the data used for pre-training. We evaluate the generalization on external datasets via two common practices to utilize the foundation model: (1) in-domain fine-tuning on separated datasets and tasks, and (2) fully external validation of the disease detection models without any site-specific fine-tuning.
@@
For in-domain fine-tuning, the foundation model is fine-tuned on each external dataset’s training set and validated on held-out sets from the same source. The bottom four rows in Fig. 1e report the tasks-level performances on NYU Long Island and RSNA datasets. The fine-tuned model yields a macro-AUC of 0.904 across the 10 tasks on NYU Long Island dataset and a macro-AUC of 0.923 for five types of hemorrhages on the RSNA dataset. In comparison, the scratch model results in macro-AUC scores of 0.748 and 0.824, respectively. Moreover, the foundation model also significantly outperforms Merlin, as shown in Supplementary Fig. 5. The superior performances on external datasets indicate the generalizability of the foundation model. Note that the limited data size of CQ500 forbids training a effective deep learning model from scratch, reinforcing the importance of the foundation model in label efficiency, which is further studied in Section "Label efficiency". Interestingly, when comparing performances across different datasets, Fig. 1e demonstrates that the AUCs of the in-domain fine-tuned model on the external dataset even exceed the AUCs achieved on the internal dataset. For instance, the fine-tuned models consistently obtained AUCs greater than 0.90 in all the hemorrhage detection tasks on RSNA dataset, surpassing the AUCs on NYU Langone data. This may be attributed to the higher label quality in radiologist-reviewed datasets, for which label noise may be better controlled by comparison to EHR-derived labels.
@@
In the full external validation without any site-specific fine-tuning (illustrated in Fig. 1c), we evaluated classification models fine-tuned on the NYU Langone training set, as-is, on the held-out validation sets from each external dataset. Fig. 1f compares performance between external validation and in-domain fine-tuning. Results show that, for the NYU Long Island and RSNA datasets —where the training set used for fine-tuning includes a sufficient number of high-quality labeled samples— in-domain fine-tuning does enhance the model performance. However, on the CQ500 dataset, with only 1,120 training samples, the in-domain fine-tuned model performs worse than the model transferred from NYU Langone, especially for EDH and SDH, which have a greater class imbalance. These comparisons highlight two typical use cases for foundation models depending on the availability of labeled data for fine-tuning. Additionally, comparing the first row of Fig. 1e and external validation in Fig. 1f, the fine-tuned model on NYU Langone achieves similar AUC values on both internal and external datasets, indicating robust generalizability to external data.
@@
Another key advantage of the foundation model is its ability to facilitate transfer learning and fine-tuning tasks with minimal labeled data. For example, as shown in Fig. 1c, the CQ500 dataset contains only 1,585 scans. Despite the small dataset size, fine-tuning our foundation model on CQ500 achieves promising results, with an AUC of 0.863.
To systematically evaluate the label efficiency of our foundation model, we also assess the generalization capabilities of models on new tasks given a limited number of examples within the paradigm of few-shot learning, where only $K$ positive and negative samples each are used for training in each task. Since the quality of few-shot learning is largely determined by the sampled $K$-shots training data, we re-sampled and re-trained the model 5 times for calculating means and confidence intervals. As expected, Fig. 2 shows that performance improves as more data is used for training, with narrower confidence intervals. Surprisingly, even with a small number of examples (e.g., 512 total, with $K=256$), the model achieves performance comparable to training with the full dataset, which contains over at least 16 times more training examples in the RSNA. Notably, for tasks like detecting IVH in the RSNA dataset, the 8-shots model achieves an AUC above 0.90, a result that rivals full-data training. These findings suggest that our foundation model has learned diverse and expressive features/representations during SSL pre-training, making it highly effective for new tasks even when trained on small labeled datasets.
@@
Scaling laws have proven effective in enhancing the performance of foundation models by increasing the size of the training dataset [39]. This phenomenon is not only observed in natural language and image domains [40, 41], but also extends to medical imaging [14, 42]. As shown in Fig. 3, scaling up the foundation model by incorporating more data during self-supervised pre-training significantly improves downstream tasks performances. We compared models pre-trained with varying proportions of the available data — $10\%$, $30\%$, and $100\%$ (full dataset), observing that larger pre-training datasets consistently led to better downstream task performance. These findings highlight the potential of leveraging more data to achieve superior results, further suggesting the value of multi-institutional collaboration and federated approaches to aggregating larger datasets to enhance model quality. Noticeably, the performance for CQ500 does not change a lot from $10\%$ to $30\%$, but $100\%$ gives a sudden performance improvement, this indicates that for smaller datasets like CQ500, scaling up the data size is crucial for learning meaningful representations.

### v2

# 3D Foundation Model for Generalizable Disease Detection in Head Computed Tomography

@@ Abstract
Head computed tomography (CT) imaging is a widely-used imaging modality with multitudes of medical indications, particularly in assessing pathology of the brain, skull, and cerebrovascular system. It is commonly the first-line imaging in neurologic emergencies given its rapidity of image acquisition, safety, cost, and ubiquity. Deep learning models may facilitate detection of a wide range of diseases. However, the scarcity of high-quality labels and annotations, particularly among less common conditions, significantly hinders the development of powerful models. To address this challenge, we introduce FM-HCT: a Foundation Model for Head CT for generalizable disease detection, trained using self-supervised learning. Our approach pre-trains a deep learning model on a large, diverse dataset of 361,663 non-contrast 3D head CT scans without the need for manual annotations, enabling the model to learn robust, generalizable features. To investigate the potential of self-supervised learning in head CT, we employed both discrimination with self-distillation and masked image modeling, and we construct our model in 3D rather than at the slice level (2D) to exploit the structure of head CT scans more comprehensively and efficiently. The pre-training phase is followed by fine-tuning on smaller, annotated downstream datasets, thereby optimizing the model for specific diagnostic tasks, such as detecting hemorrhages, tumors, and other abnormalities. The model’s downstream classification performance is evaluated using internal and three external datasets, encompassing both in-distribution (ID) and out-of-distribution (OOD) data. Our results demonstrate that the self-supervised foundation model significantly improves performance on downstream diagnostic tasks compared to models trained from scratch and previous 3D CT foundation models on scarce annotated datasets. Furthermore, the model maintains strong generalization across different datasets, indicating its potential for broad clinical applicability. This work highlights the effectiveness of self-supervised learning in medical imaging and sets a new benchmark for head CT image analysis in 3D, enabling broader use of artificial intelligence for head CT-based diagnosis.
@@ Introduction
Head computed tomography (CT) is often the first step in diagnosing a wide range of neurological disorders, including head trauma, hemorrhages, hydrocephalus, and malignancies. Head CT scans are faster, more accessible, and generally less expensive than magnetic resonance imaging (MRIs), making them ideal for emergencies like traumatic brain injury (TBI) or suspected stroke. They are also effective in detecting bone fractures, or neurovascular pathologies such as arterial venous malformations. Despite its widespread use, CT lacks the contrast resolution and hence the sensitivity for many disorders dependent upon diagnosis by MRI, thus MRI is the imaging modality of choice for many neurologic diseases. MRI, however, is more costly, risks potential heating or displacement of indwelling implants, and suffers generally slower acquisition times, increasing patient discomfort and risking non-diagnostic examinations due to its greater sensitivity to motion-related artifacts. It is also more expensive than CT and is contraindicated in specific patients. Access to MRI is a major challenge in resource-limited countries. The timely and arduous determination of certain pathologies can delay appropriate medical and surgical treatment for patients.
There is significant potential to harness artificial intelligence (AI) algorithms to enhance the diagnostic and early detection capabilities of head CT, providing critical support in clinical decision-making and improving patient outcomes. Early and accurate diagnosis can potentially lead to more effective treatments, reduce complications, and improve patient survival.
@@
Current research on AI-driven diagnosis using head CT is limited due to both lack of data availability and the complexity of model architectures. Although datasets such as RSNA [1] and CQ500 [2] provide publicly available head CT data, they remain small (RSNA includes approximately 10K samples and CQ500 approximately 1K), and their primary focus is on hemorrhage detection, which restricts broader applicability as a credible path to clinical decision support. Moreover, many existing models are designed with highly task-specific architectures that may not generalize well to diverse clinical applications. These models typically apply 2D convolution neural networks (CNN) to sequentially process 3D volumes slice-by-slice under the supervision of slice-level labels [3, 2, 4]. Slice-level labels are often expensive to acquire, and models trained on 2D slices often struggle to generalize to conditions like neurodegenerative diseases, where slice-level labels are not easily defined. Developing models that can harness the information embedded within the 3D structure of CT images while requiring minimal slice-level labeled data can thus expand the impact of such approaches. To address these challenges, we have developed FM-HCT: a Foundation Model for Head CT, and demonstrated robust performance across multiple tasks and datasets, which highlights our model’s potential for broad clinical applicability.
@@
In this work, we introduce FM-HCT, a 3D foundation model for head CT scans, developed using self-supervised learning (SSL). While SSL has shown success in natural images [21, 22, 23, 24, 25, 26, 6, 7, 27, 28] and in medical imaging [29, 30, 12, 14, 31, 32, 13, 33, 34, 35, 36], training a robust 3D CT volume encoder presents distinct challenges, such as selecting appropriate pretext tasks, managing spatial normalization, and addressing high computational demands. To overcome these challenges, we developed a standardized pipeline that normalizes head CT scans from various protocols, producing consistent input for our foundation model. For pre-training, we adapted two SSL frameworks—self-distillation, inspired by DINO [24], and masked prediction, inspired by MAE [26]. These methods were tailored to train a volumetric encoder based on a customized vision Transformer (ViT)[37]. The full details of our design and choices are described in Section “Method”. pre-training was conducted on a large-scale dataset comprising 361,663 head CT scans from a major clinical institution.
@@
To evaluate the foundation model, we systematically assessed its performance and generalizability across 10 downstream disease detection tasks using diverse internal and external datasets, as illustrated in Fig.˜1. Beyond commonly studied hemorrhages, our evaluation includes crucial yet less-explored tasks in head CT, such as identifying brain tumors, Alzheimer’s disease and related dementia (ADRD), edema, and hydrocephalus (HCP). For each downstream task, the foundation model was fine-tuned using task-specific labels. Given the scarcity of expert-annotated public datasets for these conditions, we leverage electronic health records (EHR) to acquire labels of each task. While EHRs may include missing data and suffer potential label-noise, they remain a valuable and practical source for large-scale patient status labeling that can be used to evaluate the performance of the foundation model. In order to assess the label quality, we provide label sensitivity analysis by comparing 1000 manually labeled cases by one radiologist in Supplementary LABEL:tab:label_noise.
@@
Our results reveal substantial performance improvements enabled by our foundation model. Downstream models initialized with the pre-trained weights of foundation model achieve a $16.07\%$ improvement in macro-AUC over models trained from scratch with random initialization on internal NYU Langone data, and $20.86\%$ and $12.01\%$ improvements on external datasets from NYU Long Island (previously a separate hospital) and RSNA, respectively ($P<0.001$ for all comparisons). These findings underscore the potential of our foundation model to advance AI-based interpretation of head CT scans, supporting more accurate diagnosis and early disease detection. Furthermore, as described in Section “Results”, we demonstrate the model’s capabilities in out-of-distribution generalization (Fig.˜1), few-shot learning (Fig.˜2), and scalability (Fig.˜4), highlighting the method’s potential in scenarios with limited annotated fine-tuning data, or scenarios such as federated learning which provide access to orders of magnitude more data. Overall, the experimental results on multiple datasets and tasks underscore the generalizability, adaptability and effectiveness of the model, and pave the way for significant impact in real-world clinical applications.
@@
Figure 1: Overview of the study - the approach to developing a foundation model for head CT and its performance in disease detection tasks. $n$ in the Figure refers to number of samples for each dataset. a, Collection of training data and pretraining of the foundation model. b, Query disease labels associated with head CT scans for downstream tasks. c, Evaluation design of the foundation model using both internal and external datasets. d, Application of the foundation model to various downstream disease detection tasks. e-g, The performance comparison among different scenarios: (e) training with and without foundation model, (f) comparison of our CT foundation model vs. other CT foundation models, (g) in-domain finetuning and external validation in transfer learning.
@@ Results
The key aim of the foundation model is to develop a single model that improves performance on a wide range of downstream tasks of detecting recognizable abnormalities from head CT scans. To evaluate the capability of the foundation model, we train classification models for multiple disease detection tasks by fine-tuning the foundation model separately per disease, and assessing the fine-tuned model’s performance on held-out validation and external data sets. The selected downstream tasks include detecting various types of hemorrhages (intraparenchymal hemorrhage (IPH), intraventricular hemorrhage (IVH), subdural hemorrhage (SDH), epidural hemorrhage (EDH), subarachnoid hemorrhage (SAH), and intracranial hemorrhage (ICH)), brain tumors, hydrocephalus (HCP), edema, and Alzheimer’s diseases and related dementia (ADRD). Fig.˜1a,b,c show the overview of our pre-training framework and included data, EHR-matching, and datasets used in pre-training, in-domain fine-tuning, and external validation. Overall N=361,663 scans were used during pre-training, and four distinct datasets from different sources were used for various forms of validation (NYU Langone N=26,487; NYU Long Island N=2,202; RSNA N=1,058; and CQ500 N=236). NYU Langone is a hospital system comprised of multiple geographically distinct hospitals including two Level 1 Trauma Centers and three Comprehensive Stroke Centers. NYU Long Island, a Level 1 Trauma Center/Comprehensive Stroke Center, is treated as an external dataset for the purposes of this study.
@@
The first two rows of Fig.˜1e report the task-specific AUCs for Vision Transformer (ViT) classifiers trained from scratch with random initialization, namely scratch, versus those fine-tuned from the foundation model, namely fine-tuned on NYU Langone data. The fine-tuned models consistently outperform the scratch model across all 10 disease detection tasks, achieving a macro-AUC of 0.852 — a $16.07\%$ relative increase over the scratch model’s 0.734 ($P<0.001$). Additionally, in Fig.˜1f and Supplementary Fig.˜5 we compared the foundation model with three other foundation model for 3D CT scans — Merlin [20], Google’s CT Foundation [17] model and CT-FM [38]. Across four compared datasets illustrated in Supplementary Fig.˜5, Merlin outperforms the scratch model with a macro-AUC relative improvement of $8.07\%$ and $27.50\%$ on Average Precision (AP) while falling short compared to our foundation model with $13.05\%$ relative lower macro-AUC ($P<0.001$) and $27.50\%$ on AP ($P<0.001$). Although Merlin is not directly comparable to our foundation model as it was pre-trained on abdominal CT, it still provides a valuable baseline. To complement the shortcoming that Merlin is a baseline not pre-trained on Head CT, we compare our model with CT-FM [38], a foundation model pretrained on 148,000 diverse CT scans, including head CT. Our model demonstrates an macro-AUC relative improvement of $9.56\%$ ($P<0.001$) and $44.60\%$ on AP ($P<0.001$) as shown in Supplementary Fig.˜5, demonstrating the effectiveness and scalability of our approach. Additionally, we compare our model to Google CT Foundation model with linear probing, because trainable weights for end-to-end fine-tuning are not provided for this model. We consistently observe improved model performance across the board (in Fig.˜1f and Supplementary Fig.˜7). These findings demonstrate that despite the progress in general domain multimodal models, specialized foundation model pre-trained on head CT data still significantly enhance the understanding of brain diseases.
@@
To assess our foundation model’s generalization to out-of-distribution data, we compiled three external datasets from multiple institutions and sources: NYU Long Island, RSNA [1], and CQ500 [2], as shown in Fig.˜1c (NYU Langone and NYU Long Island are geographically separate and distinct institutions within the broader health system). The data in these external datasets has a different distribution than the data used for pre-training. We evaluate the generalization on external datasets via two common practices to utilize the foundation model: (1) in-domain fine-tuning on separated datasets and tasks, and (2) fully external validation of the disease detection models without any site-specific fine-tuning.
@@
For in-domain fine-tuning, the foundation model is fine-tuned on each external dataset’s training set and validated on held-out sets from the same source. The bottom four rows in Fig.˜1e report the tasks-level performances on NYU Long Island and RSNA datasets. The fine-tuned model yields a macro-AUC of 0.904 across the 10 tasks on NYU Long Island dataset and a macro-AUC of 0.923 for five types of hemorrhages on the RSNA dataset. In comparison, the scratch model results in macro-AUC scores of 0.748 and 0.824, respectively. Moreover, the foundation model also significantly outperforms Merlin, as shown in Supplementary Fig.˜5. The superior performances on external datasets indicate the generalizability of the foundation model. Note that the limited data size of CQ500 forbids training a effective deep learning model from scratch, reinforcing the importance of the foundation model in label efficiency, which is further studied in Section "Label efficiency". Interestingly, when comparing performances across different datasets, Fig.˜1e demonstrates that the AUCs of the in-domain fine-tuned model on the external dataset even exceed the AUCs achieved on the internal dataset. For instance, the fine-tuned models consistently obtained AUCs greater than 0.90 in all the hemorrhage detection tasks on RSNA dataset, surpassing the AUCs on NYU Langone data. This may be attributed to the higher label quality in radiologist-reviewed datasets, for which label noise may be better controlled by comparison to EHR-derived labels.
@@
In the full external validation without any site-specific fine-tuning (illustrated in Fig.˜1c), we evaluated classification models fine-tuned on the NYU Langone training set, as-is, on the held-out validation sets from each external dataset. Fig.˜1f compares performance between external validation and in-domain fine-tuning. Results show that, for the NYU Long Island and RSNA datasets —where the training set used for fine-tuning includes a sufficient number of high-quality labeled samples— in-domain fine-tuning does enhance the model performance. However, on the CQ500 dataset, with only 1,120 training samples, the in-domain fine-tuned model performs worse than the model transferred from NYU Langone, especially for EDH and SDH, which have a greater class imbalance. These comparisons highlight two typical use cases for foundation models depending on the availability of labeled data for fine-tuning. Additionally, comparing the first row of Fig.˜1e and external validation in Fig.˜1f, the fine-tuned model on NYU Langone achieves similar AUC values on both internal and external datasets, indicating robust generalizability to external data.
In addition to in-domain fine-tuning model performance comparison present in Supplementary Fig.˜5. Out-of-domain generalizability of our model is assessed in Fig.˜1g, where our model is first fine-tuned on in-domain finetuning dataset and then evaluated on external datasets (NYU Longisland, RSNA and CQ500). The result shows our model can demonstrate comparable performance on external dataset.
Volume to volume hemorrhage sub-types retrieval performance
To further rigorously evaluate the representation quality of our pre-trained foundation model, we conducted a volume-to-volume hemorrhage subtyping retrieval study on RSNA and CQ500, comparing against Google CT, CT-FM, and Merlin (Fig.˜3).
Retrieval performance was assessed using mean Average Precision (Retrieval mAP), which is conceptually distinct from the Average Precision (AP) metric used earlier in the paper for classification tasks. In the retrieval setting, mAP is computed as the mean of all individual query-level AP scores, where each AP score is defined as the average precision at the ranks where relevant items are retrieved, i.e., at positions $K=\{k_{1},k_{2},...,k_{R}\}$. In our study, we report AP at ranks $K=\{1,5,10\}$. A more detailed formulation of retrieval mAP can be found in the “Method” Section.
Our model demonstrates substantial improvements over CT-FM and Merlin, and achieves notable relative gains of $9.99\%$ on CQ500 and $2.21\%$ on RSNA compared with Google CT, averaged across hemorrhage subtypes. Additional retrieval methodological details are provided in “Method” Section, and a comprehensive evaluation with Precision@K@K is reported in the Supplementary Fig.˜10.
          RSNA            CQ500
Refer to caption Refer to caption
Figure 3: Volume-to-Volume Retrieval Performance Comparison The plot presents mean Average-Precision (Retrieval mAP) for volume-to-volume retrieval with hemorrhage sub-types retrieval on RSNA and CQ500. All vs. all image retrieval is performed in this study, where every image in the dataset is used as a query once, and the gallery (the search space) is the entire dataset itself. Additional methodological details are provided in “Method” Section. Additional evaluation on Precision@K@K is present in Supplementary Fig.˜10. In the plot, we show that our model shows better retrieval performance compared alternative models in majority of cases.
@@
Another key advantage of the foundation model is its ability to facilitate transfer learning and fine-tuning tasks with minimal labeled data. For example, as shown in Fig.˜1c, the CQ500 dataset contains only 1,585 scans. Despite the small dataset size, fine-tuning our foundation model on CQ500 achieves promising results, with an AUC of 0.863.
To systematically evaluate the label efficiency of our foundation model, we also assess the generalization capabilities of models on new tasks given a limited number of examples within the paradigm of few-shot learning, where only $K$ positive and negative samples each are used for training in each task. Since the quality of few-shot learning is largely determined by the sampled $K$-shots training data, we re-sampled and re-trained the model 5 times for calculating means and confidence intervals. As expected, Fig.˜2 shows that performance improves as more data is used for training, with narrower confidence intervals. Surprisingly, even with a small number of examples (e.g., 512 total, with $K=256$), the model achieves performance comparable to training with the full dataset, which contains over at least 16 times more training examples in the RSNA. Notably, for tasks like detecting IVH in the RSNA dataset, the 8-shots model achieves an AUC above 0.90, a result that rivals full-data training. These findings suggest that our foundation model has learned diverse and expressive features/representations during SSL pre-training, making it highly effective for new tasks even when trained on small labeled datasets.
We additionally show few-shot model performance on CT-FM in Supplementary Figs.˜11 and 12, where it shows lower few-shot capability in comparison to our model.
@@
Comparison to alternative modeling choices
In order to further verify the effectiveness of our proposed method on modeling 3D CT scans, we additionally compare to two alternative modeling choices — 1) multiple instance learning (ABMIL [39]) with a state-of-the-art 2D foundation model DINOv3 [40], 2) mean pooling with DINOv3, 3) modeling 3D CT Scans as video with a state-of-the-art video foundation model VJEPA2 [41]. The detailed model fine-tuning comparison across four datasets is provided in Supplementary Fig.˜8. Since both DINOv3 and VJEPA2 support dynamic resolutions, we keep the input size to be the same as our model for fair comparison ($96\times 96\times 96$). The result shows that our model perform best for majority of tasks with $2.57\%$ relative improvement on AUROC and $15.54\%$ on AP against DINOv3 with ABMIL, $9,26\%$ relative improvement on AUROC and $46.57\%$ on AP against DINOv3 with mean pooling, and $6.26\%$ relative improvement on AUROC and $26.25\%$ on AP against VJEPA2. Furthermore, from model efficiency analysis in Supplementary Fig.˜9, we show that our modeling method has significant advantage over other benchmarked architectures in term of both model throughput and memory cost.
@@
Scaling laws have proven effective in enhancing the performance of foundation models by increasing the size of the training dataset [42]. This phenomenon is not only observed in natural language and image domains [43, 44], but also extends to medical imaging [14, 45]. As shown in Fig.˜4, scaling up the foundation model by incorporating more data during self-supervised pre-training significantly improves downstream tasks performances. We compared models pre-trained with varying proportions of the available data — $10\%$, $30\%$, and $100\%$ (full dataset), observing that larger pre-training datasets consistently led to better downstream task performance. These findings highlight the potential of leveraging more data to achieve superior results, further suggesting the value of multi-institutional collaboration and federated approaches to aggregating larger datasets to enhance model quality. Noticeably, the performance for CQ500 does not change a lot from $10\%$ to $30\%$, but $100\%$ gives a sudden performance improvement, this indicates that for smaller datasets like CQ500, scaling up the data size is crucial for learning meaningful representations.